OpenStudy (anonymous):

how do i fin the y intercept for the function y=0.5(1/4)^x ?????MEDAL

OpenStudy (anonymous):

replace \(x\) by \(0\) that is how you find the y intercept of any function it is always \(f(0)\)

OpenStudy (misty1212):

HI!!

OpenStudy (here_to_help15):

Hello MISTY :D

OpenStudy (misty1212):

\[\left(\frac{1}{4}\right)^0=1\] just like any number to the power of zero is one

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

what do you like how does that look written out??? HI(:

OpenStudy (anonymous):

okay so the y intercept is 1?

OpenStudy (misty1212):

\[y=0.5\left(\frac{1}{4}\right)^0=0.5\times 1=0.5\] that is written out

OpenStudy (anonymous):

i thought 0.5 is the slope?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

just someone okay i get the process of how you did that but i just want to know 100% that it is 0.5 even though the slope is the same @misty1212

OpenStudy (er.mohd.amir):

intercept is zero y=mx+c here y-intercept is c compare your eq with this

OpenStudy (anonymous):

okay so its actually 0 in every case that ^x is in a function

OpenStudy (anonymous):

this is an exponential function, it is not a line, it does not have a slope

OpenStudy (anonymous):

ugh thank you i was so confused it looked like mx+b and it was asking for a y intercept thank you!!!! but this is exponential decay and y intercept is no solution